OpenStudy (anonymous):

solve for g. s=4g+4w

OpenStudy (anonymous):

s = 4g + 4w 4g = s - 4w g = s/4 - w Just have to work with a bit of equation simplification and moving things around. :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would i find the answer?

OpenStudy (anonymous):

g = s/4 - w is the answer. Unless you know what the values of s and w are. Sorry, I have to go now. Hope I helped!

OpenStudy (texaschic101):

I totally agree....g = s/4 - w

OpenStudy (anonymous):

it says its wrong

OpenStudy (texaschic101):

what about this.. s = 4g + 4w --- subtract 4w from both sides s - 4w = 4g --- divide by 4 (s - 4w)/4 = g